Raymond Mining Corporation has 9.7 million shares of common stock outstanding, 410,000 shares of 4% $100 par value preferred stock outstanding, and 167,000 7.50% semiannual bonds outstanding, par value $1,000 each. The common stock currently sells for $45 per share and has a beta of 1.35, the preferred stock currently sells for $94 per share, and the bonds have 10 years to maturity and sell for 116% of par. The market risk premium is 8.5%, T-bills are yielding 5%, and Raymond Mining’s tax is 35%.

a. What is the firm’s market value capital structure?

Market value
Debt
Equity
Preferred stock

b. If Raymond Mining is evaluating a new investment project that has the same risk as the firm’s typical project, what rate should the firm use to discount the project’s cash flows? (Do not round intermediate calculations. Enter your answer as a percentage rounded to 3 decimal places.)

Discount rate          
Answer #1
Solution:
Given:
No of Common stock 9.7 Million shares i.e 97000000 shares
No of Preferred stock 410000 shares
No of Semiannual bonds 167000 Bonds
Current Market value of common stock $45 per share 97000000
Current Market value of Preferred stock $94 per share
Current Market value of Bonds 116% of par value i.e 116%*$1000
a)
Quantity i.e Number Market value Per share/Bond Total Market value Weights Working
a b c = a*b
Debt 167000 $1160 ($1000*116%) $193720000 0.04213814 $193720000/$4597260000
Equity 97000000 $45 $4365000000 0.9494786 $4365000000/$4597260000
Preferred stock 410000 $94 $38540000 0.00838325 $38540000/$4597260000
Total Market value $4597260000 1
b)
As new investment has same risk , we would use weighted average cost of capital (WACC )to discount the project cash flows
WACC = (Cost of Debt * weight of Debt )+(Cost of Equity *Weight of Equity )+(Cost of Preferred stock *Weight of preferred stock)
Calculation of cost of debt :
To calculate Yield to maturity debt we will use RATE Function in excel:
Given:
Current price of bond i.e PV $1160
Par value i.e FV $1000
Coupon amount i.e PMT $37.5 ($1000*7.5%/2) As bonds are semiannual , therefore divided by 2
Term of Bond i.e NPER 10 Years i.e 10*2 = 20 periods, as bonds are semiannual
RATE(20,$37.5,-$1160,$1000,0)
2.703785%
YTM on bond = 2.703785*2 0.054075703
YTM on bond = 5.407570283% 5.407570283
Cost of debt = Yield on bond *(1-Tax rate)
5.407570283% *(1-0.35)
5.407570283% *(1-0.35)
0.035149207
3.514920684
Cost of debt = 3.514920684%
Calculation of cost of Equity :
Cost of Equity = Risk free rate (T-Bill rate)+Beta*(Risk Premium)
5%+(1.35*8.5%)
0.16475
16.475
Cost of Equity = 16.475%
Calculation of cost of Preferred stock:
Cost of Preferred stock = Preferred dividend/Current price of Preferred stock
Dividend = Par value *rate of dividend
$100*4% =$4
Cost of Preferred stock = $4/$94
0.042553191
4.255319149
Cost of Preferred stock = 4.25531914893617
Calculation of weighted average cost of capital
Components Weight Cost Weighted average cost
a b c = a*b
Debt 0.042138143 3.514920684 0.148112231
Equity 0.949478602 16.475 15.64265998
Preferred stock 0.008383254 4.255319149 0.035673423
15.82644563
The firm should discount the project at 15.826%